OpenStudy (anonymous):

The differential equation dy dx equals the quotient of the quantity y minus 2 and y plus 1 produces a slope field with horizontal tangents at y = 2 produces a slope field with vertical tangents at y = −1 produces a slope field with columns of parallel segments

OpenStudy (anonymous):

https://gyazo.com/40b80568af9602738d4d7c64db1bb2e1

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

I was thinking C since the denominator would be bigger.

zepdrix (zepdrix):

\[\large\rm y'(y)=\frac{y-2}{y+1}\]At y=2,\[\large\rm y'(2)=\frac{2-2}{2+1}\]Hmm, so what is this telling us?\[\large\rm y=2, \qquad\qquad y'=0\]

zepdrix (zepdrix):

Oh I didn't look at the options yet hehe

OpenStudy (anonymous):

Horizontal line

zepdrix (zepdrix):

So first line is definitely true, cool.

OpenStudy (anonymous):

10/10 IGN

zepdrix (zepdrix):

Looks like the second one is true also, ya? Bottom blows up, corresponding to "infinite" slope, asymptotic behavior.

zepdrix (zepdrix):

I didn't even look at the third option yet :P Just seems like we land on C by checking the first two automatically lol

zepdrix (zepdrix):

Clearly III can't be true if I and II are true, right? If we have vertical tangents, and we also have horizontal tangents in other places, then we don't have parallel tangent lines.

OpenStudy (anonymous):

I was thinking about it in a completely different way.

OpenStudy (anonymous):

Then I noticed it was asking about the slope fields.

OpenStudy (anonymous):

In all three where as I misread I and II and eliminated them using some thought (yes I thought wrong ;p) thanks